Lawn Care, Inc., has sales of $367,400, costs of $183,600, depreciation of $48,600, interest of $39,200, and a tax rate of 25 percent. The firm has total assets of $422,100, long-term debt of $102,000, net fixed assets of $264,500, and net working capital of $22,300. What is the return on equity?

Multiple Choice

  • 24.26 percent

  • 15.38 percent

  • 38.96 percent

  • 29.96 percent

  • 17.06 percent
Answer #1
Sales 367400
Costs (183600)
Depreciation (48600)
EBIT 135200
Less:interest expense (39200)
EBT 96000
Less:taxes(96000*25%) (24000)
Net income 72000

Total assets=Fixed assets+Current assets

Hence current assets=422100-264500

=$157600

Net working capital=Current assets-Current liabilities

Hence current liabilities=(157600-22300)=$135300

Total assets=Total liabilities+Equity

422100=(135300+102000)+equity

equity=422100-(135300+102000)

=$184800

ROE=Net income/equity

=72000/184800

=38.96%(Approx).